I need 30% of a 60% solution. I currently have a 40% solution, and a 70% solution. How many liters of each do I need to make 30 liters of 60% solution? Muchas Gracias
Answer by stanbon(75887)

I need 30 liters of a 60% solution. I currently have a 40% solution, and a 70% solution. How many liters of each do I need to make 30 liters of 60% solution?
------------------
Equation:
active + active = active
0.40x + 0.70(30-x) = 0.60*30
------
40x + 70*30 - 70x = 60*30
-30x = -10*30
-----
x = 10 liters (amt. of 40% solution needed)
------
60-x = 50 liters (amt. of 70% solution needed)
==============
Cheers,
Stan H.
===================